OpenStudy (anonymous):

A container has a large cylindrical lower part with a long thin cylindrical neck. The lower part of the container holds 12.5 m3 of water and the surface area of the bottom of the container is 4.61 m2. The height of the lower part of the container is h1 = 2.90 m and the neck contains a column of water h2 = 7.80 m high. The total volume of the column of water in the neck is 0.200 m3. What is the magnitude of the force exerted by the water on the bottom of the container?

OpenStudy (ujjwal):

pressure at bottom= density * g * h =1000 * 9.8 * 10.7 =104860 N/m2 Now force = pressure * area = 104860 * 4.61= 483404.6